# Reducing Numerical Precision Requirements in Quantum Chemistry Calculations

### v1

@@ Abstract
I Introduction
In recent years, progress in the field of Artificial Intelligence has lead to an increase in demand for computing resources to perform deep learning [1]. This has significant implications for developments in computational quantum chemistry — simulation software must be written to coexist on AI-centric software and hardware ecosystems. Of particular importance will be the ability to run quantum chemistry packages on low precision hardware such as NVIDIA’s Tensor Cores or Google’s TPUs (see, for example, Ref. [2] and Ref. [3] respectively for their application to scientific problems). While targeting low precision hardware will increase the challenge of developing simulation software in the short term, it also presents new opportunities for co-designing specialized hardware optimal for solving scientific problems.
In this paper, we will systematically explore the floating point precision requirements necessary for a representative quantum chemistry kernel: calculation of the single particle density matrix represented in a linear combination of atomic orbitals (LCAO) basis set. We will find that double precision calculations provide an unnecessarily high level of precision. We will then propose the use of the error-free transformation for matrix multiplication developed by Ozaki and coworkers [4] (i.e. the Ozaki scheme) in combination with density matrix purification [5] to exploit low precision hardware while obtaining the required precision (and no more).
II Background
@@
II.1 Floating Point Representations
@@
II.2 Low Precision Quantum Chemistry Calculations
@@
There has been significant research on low precision computing for computational quantum chemistry, however it has primarily focused on single precision vs. double precision. Single precision for the analytic calculation of two-electron ERIs in a Gaussian basis set has been demonstrated to be an effective strategy to exploit GPUs [6, 7, 8, 9]. Single precision can also be employed to accelerate semi-numerical methods [10]. The single precision strategy has recently been applied to Slater type orbitals as well [11]. In materials science, single precision was shown to be a promising strategy to accelerate the computation of exact-exchange in codes using a planewave basis set [12]. Single precision has also been employed to speed up the iterative eigenvalue solvers of materials codes [13, 14, 15, 16]. Single precision calculations are particularly promising for many body perturbation theory methods. Early work demonstrated the reduced precision requirements of MP2 calculations [17, 18], including when using the RI technique [19]. Single precision can also be applied to Coupled Cluster [20, 21], including time dependent variants [22]. Other targets of single precision optimizations include DMRG [23], quantum transport calculations [24], and GW [25].
@@
A challenge the community has faced for developing low precision software has been to predict how relevant such optimizations will be to future architectures. For example, Yasuda showed that evaluation of the exchange and correlation functional could be accelerated using single precision [26]; however, a recent work [27] explicitly rejected this strategy noting the gap in performance between double and single precision on GPUs has been closed. The earlier mentioned work on semi-numerical calculations [10] justified its strategy by targeting lower-cost “gaming GPUs”, where single precision continues to have higher performance. With recent developments in Artificial Intelligence, the pendulum has swung back towards the relevance of low (and more exotic) precision hardware. It is now crucial for the quantum chemistry community to establish clear precision requirements for their simulations. We note that even if double precision capable hardware remains the standard, studying the effects of low precision will still be potentially useful for reducing data transfer costs.
II.3 Density Matrix Purification
@@
In mean field quantum chemistry calculations, such as Kohn-Sham Density Functional Theory [28, 29], we need to compute the single particle density matrix from a given Hamiltonian. Limiting ourselves to the spin-restricted case, we expand the orbitals in some set of $M$ basis functions:
@@
In most implementations based on LCAOs, equation 4 is solved by invoking a dense eigenvalue solver, such as the ones available in LAPACK or ScaLAPACK. Unfortunately, these calculations have a computational cost that scales with the third power of the number of basis functions. To reduce this cost for application to large systems, many “diagonalization free” methods have been proposed [30]. One class of “diagonalization free” methods is based on the purification algorithm first proposed by McWeeny [31], which iteratively computes $K$ using the following recurrence relation:
@@
The power of such an approach is that the core computational kernel is matrix - matrix multiplication, which can readily exploit the underlying sparsity of $H$ and $K$ that exist for large insulating systems [32]. A number of different purification methods exist (see, for example, the methods implemented in the NTPoly library [33]); each employs different forms and orders of polynomials during the iterations. The second order trace correcting method of Niklasson [34] has the benefit of only needing to compute the square of a matrix, a point we will return to in Sec. IV.5.
@@
Density matrix purification is not only useful for the case of extremely large systems where such sparsity exists. Since the purification algorithm has matrix - matrix multiplication as a bottleneck, it has the benefit of scaling better on supercomputers than eigenvalue solvers do. When developing a Hartree-Fock code for the Tianhe-2 supercomputer, Chow and coworkers utilized purification to substantially improved scalability [35, 36]. Finkelsein and coworkers recently demonstrated how an algorithm similar to purification implemented on a GPU could outperform dense eigenvalue solvers even on a single GPU [37].
@@
Pederson and coworkers [38] similarly have proposed using dense purification as a means of exploiting a cluster of Google TPUs. In their work, they use a mixed-precision scheme where early SCF iterations are performed in single precision and the final iterations in a software emulated double precision. Around the same time, Finkelstein and coworkers performed a series of studies using the Tensor Cores available on NVIDIA GPUs [39, 40]. They targeted single precision accuracy by taking advantage of the Tensor Core’s ability to accumulate in single precision and employing the Markidis scheme [2]. The work was further extended to density functional perturbation theory [41]. In our work here, we will go beyond these earlier studies to achieve the higher precision required for general application.
II.4 Ozaki Scheme
@@
The Ozaki scheme [4] performs an error-free transformation of computing the product of two matrices into a summation of several matrix multiplications that can be performed without rounding error. Several implementations of the Ozaki scheme exist both to target higher than double precision [42, 43] and for using low precision units like Tensor Cores [44, 45]. Remarkably, an implementation based on the INT8 Tensor Cores of an RTX A6000 GPU could outperform double precision cuBLAS by >4×>4$\times$> 4 × without loss of accuracy. On future architectures where low precision arithmetic units further dominate FP64, this scheme would be even more potent.
@@
The details of the Ozaki scheme have been presented in several previous publications, so we only briefly review the concept here (Fig. 2). The scheme begins by splitting the input matrices into $S$ split matrices (of the same size as the original). This is done through a series of rounding and bit shifting operations so that each matrix can be represented exactly in the low precision representation (we include the improved version introduced later by Minamihata et al. [46], see equation 3 in the paper of Mukunoki et al. [47] for details). A further scaling operation is applied to maintain the exponent’s range [44] (we note that in the scaling method of Mukunoki, only error free terms are considered, but in our implementation we extend the scaling to all terms). We then compute the product of pairs $A_{i}B_{j}$, where $i+j\leq S+1$ as well as a set of remainder terms. Finally, the resulting matrices are summed up in the original precision. The accuracy of the final result depends on the number of splits; if both matrices are split $S$ times we require $2S$ times as much memory and $S(S+1)/2$ as many multiplications.
@@
In Fig. 3, we show some example calculations of the multiplication of two random matrices (FP64 elements distributed between [0, 1]) using different precisions and splits. Accuracy is measured as the Frobenius norm of the difference between the double precision reference result and the Ozaki scheme result. From this data, we see one crucial point about the Ozaki scheme: the floating point precision for accumulation determines the overall precision. Hence the Ozaki-scheme is particularly effective at taking advantage of hardware like NVIDIA’s Tensor Cores.
III Implementation Details
@@
For this study, we created two different implementations of density matrix purification: one based on NVIDIA’s CUDA API for use on an NVIDIA RTX A6000 GPU and another using GNU MPFR to emulate low precision operations in software. MPFR will be particularly helpful to this study as it allows us to investigate arbitrarily defined floating point numbers [48]. However, one significant limitation of MPFR is that it emulates an infinite exponent, though it is able to detect under/overflow. Thus we will use MPFR to explore precise precision requirements, and validate our results using the (substantially faster) CUDA implementation.
@@
As input to our purification implementation, we use Hamiltonians coming from the PySCF [49] and BigDFT [50] codes. We use HGH pseudopotentials [51] to remove the core electrons in BigDFT. BigDFT calculations are performed in the linear-scaling mode in order to produce a Hamiltonian in an LCAO basis set [52]. PySCF calculations are done with the Polarization Consistent basis sets series [53]. Calculations with BigDFT are performed with the PBE exchange-correlation functional [54] and for PySCF with B3LYP [55].
@@
For simplicity, we first transform equation 4 to the standard eigenvalue problem using the Löwdin method by diagonalization in double precision. For our tests, we implement the Trace Resetting Fourth Order (TRS4) purification method [34]. We choose this method because it gives a clearer convergence signal than lower order methods. As a convergence criteria for the purification iterations, we use a change in the electronic energy, $Tr(HP)$, of $1\times 10^{-8}$ Hartree; this in turn serves as a benchmark of whether we have achieved sufficient precision. For cases where convergence can’t be achieved (due to low precision errors), the purification algorithm halts when the previous two energy values have increased, with an absolute change of less than $1\times 10^{-2}$ Hartree (or on the 100th iteration). The source code of our implementation is available online (gitlab.com/wddawson/ozp).
IV Numerical Experiments
@@
As test cases, we will first use the Hamiltonian coming from a BigDFT calculation of a Molnupiravir molecule bound to six water molecules. We will then expand the dataset to include a selenite ion surrounded by 10 water molecules calculated with different basis sets in PySCF. We also will include water clusters of different sizes and bulk silicon computed with BigDFT. We finally analyze a large water cluster (573 molecules) computed with B3LYP/PCSEG using NTChem [56, 57] (due to the system size). The main systems used in this paper are show in Fig 4.
IV.1 Precision Requirements Sweep
@@
In order to understand the precision requirements for computing the density matrix, we perform density matrix purification using different custom floating point types implemented using MPFR on the Molnupiravir system. For this experiment, we assume infinite precision in the exponent. In Fig. 5, we plot the error of the resulting density matrix compared to a reference result computed with diagonalization in double precision. We vary the effective precision of the mantissa used for multiplication from 11 (half) to 53 (double). For each calculation, we use one of three fixed accumulation mantissa values (24, 37, 53) or the “same” precision as multiplication.
@@
Figure 5: Error in the density matrix when using purification vs. various mantissa sizes (including the implicit bit) for multiplication and accumulation applied to the Molnupiravir / BigDFT system (matrix dimension $147\times 147$). Different lines represent different values of accumulation precision and the horizontal axis represents the precision used for multiplication. The “same” line uses the same precision for multiplication and precision.
@@
From this data, we see that single precision is not sufficient to achieve a converged result. The error in the norm is $8.4\times 10^{-5}$, which corresponds to an absolute energy error of $5.2\times 10^{-6}$ Hartree. However, this failure comes from the accumulation phase, not the multiplication step. Thus, it is possible to store (and hence communicate) intermediate matrices in single precision if they are multiplied in a higher precision (a point we will return to in Sec. IV.6). We found that the magnitude of the error for any given multiplication in the iterations remains roughly the same, so there would be little benefit to adjusting the precision across iterations (for example, we tried using full double precision in the final iteration, but the error did not significantly improve). Importantly, we found that 37 bits of effective precision is sufficient for accumulation; for a single precision multiplication with 37 bits of accumulation, the error in the norm is $2.3\times 10^{-6}$ and energy is $4.2\times 10^{-8}$ Hartree. Thus, the double precision commonly used for quantum chemistry calculations essentially provides an unnecessarily high level of precision for computing the single particle density matrix.
@@
Figure 6: Error in the density matrix when using purification vs. various mantissa sizes (including the implicit bit) for multiplication and accumulation applied to the selenite / PCSEG-3 system (matrix dimension $1508\times 1508$). Different lines represent different values of accumulation precision and the horizontal axis represents the precision used for multiplication. The “same” line uses the same precision for multiplication and precision.
We now repeat this experiment using the selenite system with the PCSEG-3 basis set (Fig. 6). The pattern is very similar, with convergence again found around 37 bits of effective precision. For single precision multiplication and 37 effective bit accumulation, the error in norm is $2.9\times 10^{-5}$ and absolute error in energy $3.3\times 10^{-6}$ Hartree. Remarkably, the result with 53 effective bits for multiplication and accumulation has an error in norm of $4.9\times 10^{-6}$ and absolute error in energy of $2.0\times 10^{-10}$ Hartree. We note that this discrepancy does not exist if one uses standard double precision calculations, instead it comes from the infinite exponent of MPFR (indeed, mpfr_check_range detects such out of range errors during the calculation). Hence, it is the double precision result that contains errors. From this data, we can see that larger basis set calculations, where orbitals may be presented as a sum of small contributions from many diffuse functions, may benefit from a increase in the exponent range using a scaling scheme (such as the one discussed in Sec. II.4). Similar caution must be used when performing single precision calculations: the error in the absolute energy with standard single precision was $1.8\times 10^{-2}$ Hartree, whereas with MPFR (infinite exponent) it was $1.5\times 10^{-3}$ Hartree.
IV.2 Ozaki Scheme Application

### v2

@@ Abstract
1 Introduction
In recent years, progress in the field of Artificial Intelligence has lead to an increase in demand for computing resources to perform deep learning 1. This has significant implications for developments in computational quantum chemistry — simulation software must be written to coexist on AI-centric software and hardware ecosystems. Of particular importance will be the ability to run quantum chemistry packages on low precision hardware such as NVIDIA’s Tensor Cores or Google’s TPUs (see, for example, Ref. 2 and Ref. 3 respectively for their application to scientific problems). While targeting low precision hardware will increase the challenge of developing simulation software in the short term, it also presents new opportunities for co-designing specialized hardware optimal for solving scientific problems.
In this paper, we will systematically explore the floating point precision requirements necessary for a representative quantum chemistry kernel: calculation of the single particle density matrix represented in a linear combination of atomic orbitals (LCAO) basis set. We will find that double precision calculations provide an unnecessarily high level of precision. We will then propose the use of the error-free transformation for matrix multiplication developed by Ozaki and coworkers 4 (i.e. the Ozaki scheme) in combination with density matrix purification 5 to exploit low precision hardware while obtaining the required precision (and no more).
2 Background
@@
2.1 Floating Point Representations
@@
The Tensor Cores (TC) available on modern GPUs are one way to exploit these low precision data types. Tensor Cores perform a fused multiply add (FMA) operation on $4\times 4$ matrices, with the accumulation potentially performed in a higher precision than the input matrices. For example, NVIDIA FP16 TCs will operate on FP16 matrices and accumulate in FP32. NVIDIA’s TF32 TCs likewise take in TF32 matrices and accumulate in FP32. These specialized matrix operations can lead to substantial acceleration over standard floating point operations. In Tab. 1 we summarize the floating point performance for various precision combinations on different GPUs. We see that on an H100 SXM, using FP16 TCs instead of the FP64 TCs can potentially accelerate matrix multiplication by nearly 15x. The difference between FP64 and low precision operations is even larger when not using FP64 TCs (which are not always available). As demand for AI performance increases, we may anticipate this gap will increase further in the future.
GPU Model FP64 FP64 TC TF32 TC FP16 TC FP8 TC
NVIDIA V100 SXM2 6 7.8 125
NVIDIA A100 SXM 7 9.7 19.5 156 312 624
NVIDIA H100 SXM 8 34 67 494.5 990 1979
NVIDIA RTX 4500 ADA 9 0.619 79 159 317
NVIDIA L40S 9 1.4 183 366 733
AMD Instinct MI300X 9 81.7 163 654 1307 2615
Table 1: Performance metrics, in TFLOPS, of various GPU models using different floating point representations. For the FP64 values, we report both the performance with and without FP64 Tensor Cores. We note that the performance of the Tensor Cores on newer machines can be increased by a factor of two when sparsity can be exploited.
2.2 Low Precision Quantum Chemistry Calculations
@@
There has been significant research on low precision computing for computational quantum chemistry, however it has primarily focused on single precision vs. double precision. Single precision for the analytic calculation of two-electron repulsion integrals in a Gaussian basis set has been demonstrated to be an effective strategy to exploit GPUs 10, 11, 12, 13. Single precision can also be employed to accelerate semi-numerical methods 14. The single precision strategy has recently been applied to Slater type orbitals as well 15. In materials science, single precision was shown to be a promising strategy to accelerate the computation of exact-exchange in codes using a planewave basis set 16. Single precision has also been employed to speed up the iterative eigenvalue solvers of materials codes 17, 18, 19, 20, 21, 22. Single precision calculations are particularly promising for many body perturbation theory methods. Early work demonstrated the reduced precision requirements of MP2 calculations 23, 24, including when using the RI technique 25. Single precision can also be applied to Coupled Cluster 26, 27, including time dependent variants 28. Other targets of single precision optimizations include DMRG 29, quantum transport calculations 30, and GW 31.
@@
A challenge the community has faced for developing low precision software has been to predict how relevant such optimizations will be to future architectures. For example, Yasuda showed that evaluation of the exchange and correlation functional could be accelerated using single precision 32; however, a recent work 33 explicitly rejected this strategy noting the gap in performance between double and single precision on GPUs has been closed. The earlier mentioned work on semi-numerical calculations 14 justified its strategy by targeting lower-cost “gaming GPUs”, where single precision continues to have higher performance. With recent developments in Artificial Intelligence, the pendulum has swung back towards the relevance of low (and more exotic) precision hardware. It is now crucial for the quantum chemistry community to establish clear precision requirements for their simulations. We note that even if double precision capable hardware remains the standard, studying the effects of low precision will still be potentially useful for reducing data transfer costs.
2.3 Density Matrix Purification
@@
In mean field quantum chemistry calculations, such as Kohn-Sham Density Functional Theory 34, 35, we need to compute the single particle density matrix from a given Hamiltonian. Limiting ourselves to the spin-restricted case, we expand the orbitals in some set of $M$ basis functions:
@@
In most implementations based on LCAOs, equation 4 is solved by invoking a dense eigenvalue solver, such as the ones available in LAPACK or ScaLAPACK. Unfortunately, these calculations have a computational cost that scales with the third power of the number of basis functions. To reduce this cost for application to large systems, many “diagonalization free” methods have been proposed 36. One class of “diagonalization free” methods is based on the purification algorithm first proposed by McWeeny 37, which iteratively computes $K$ using the following recurrence relation:
@@
The power of such an approach is that the core computational kernel is matrix - matrix multiplication, which can readily exploit the underlying sparsity of $H$ and $K$ that exist for large insulating systems 38 — though whether sufficient sparsity exists to get a computational advantage depends on the system and choice of basis set. A number of different purification methods exist (see, for example, the methods implemented in the NTPoly library 39); each employs different forms and orders of polynomials during the iterations. The second order trace correcting method of Niklasson 40 has the benefit of only needing to compute the square of a matrix, a point we will return to in Sec. 4.5.
@@
Density matrix purification is not only useful for the case of extremely large systems where such sparsity exists. Since the purification algorithm has matrix - matrix multiplication as a bottleneck, it has the benefit of scaling better on supercomputers than eigenvalue solvers do. When developing a Hartree-Fock code for the Tianhe-2 supercomputer, Chow and coworkers utilized purification to substantially improved scalability 41, 42. Finkelsein and coworkers recently demonstrated how an algorithm similar to purification implemented on a GPU could outperform dense eigenvalue solvers even on a single GPU 43.
@@
Pederson and coworkers 44 similarly have proposed using dense purification as a means of exploiting a cluster of Google TPUs. In their work, they use a mixed-precision scheme where early SCF iterations are performed in single precision and the final iterations in a software emulated double precision. Around the same time, Finkelstein and coworkers performed a series of studies using the Tensor Cores available on NVIDIA GPUs 45, 46. They targeted single precision accuracy by taking advantage of the Tensor Core’s ability to accumulate in single precision and employing the Markidis scheme 2. The work was further extended to density functional perturbation theory 47 and for computing the inverse square root of the overlap matrix 48. In our work here, we will go beyond these earlier studies to achieve the higher precision required for general application.
We note that density matrix purification algorithms are only applicable to insulating systems. To address this, several different algorithms have been introduced that allow for the introduction of the electronic temperature while still having matrix multiplication as the bottleneck (see the Perspective of Aarons et al. 49 or some more recent methods 50, 51, 52). In Supplementary Information I, we perform some experiments using the Fermi Operator Expansion 53 as a representative finite temperature method to show that the findings of this paper transfer to the case of metallic systems.
2.4 Ozaki Scheme
@@
The Ozaki scheme 4 performs an error-free transformation of computing the product of two matrices into a summation of several matrix multiplications that can be performed without rounding error. Several implementations of the Ozaki scheme exist both to target higher than double precision 54, 55 and for using low precision units like Tensor Cores 56, 57. Remarkably, an implementation based on the INT8 Tensor Cores of an RTX A6000 GPU could outperform double precision cuBLAS by >4×>4$\times$> 4 × without loss of accuracy. On future architectures where low precision arithmetic units further dominate FP64, this scheme would be even more potent.
@@
The details of the Ozaki scheme have been presented in several previous publications, so we only briefly review the concept here (Fig. 2). The scheme begins by splitting the input matrices into $S$ split matrices (of the same size as the original). This is done through a series of rounding and bit shifting operations so that each matrix can be represented exactly in the low precision representation (we include the improved version introduced later by Minamihata et al. 58, see equation 3 in the paper of Mukunoki et al. 59 for details). A further scaling operation is applied to maintain the exponent’s range 56 (we note that in the scaling method of Mukunoki, only error free terms are considered, but in our implementation we extend the scaling to all terms). We then compute the product of pairs $A_{i}B_{j}$, where $i+j\leq S+1$, as well as a set of remainder terms. Finally, the resulting matrices are summed up in the original precision. The accuracy of the final result depends on the number of splits; if both matrices are split $S$ times we require $2S$ times as much memory and $S(S+1)/2$ as many multiplications.
The Ozaki scheme can be contrasted with the multiple component arithmetic approach, such as the popular double-double format 60. In the multiple component approach, each individual entry of the matrix is expressed as the sum of lower precision types. Henry et al. 61 used this approach to leverage BFLOAT16 FMA units to achieve FP32 precision. The Ozaki scheme, on the other hand, splits the matrix into the sum of several lower precision matrices. Thus, the Ozaki scheme can be thought of as a Structure of Arrays approach, whereas multiple component arithmetic is an Array of Structures approach. The Structure of Arrays approach can exploit well matrix engines like Tensor Cores since the inner operation remains unchanged.
@@
In Fig. 3, we show some example calculations of the multiplication of two random matrices (FP64 elements distributed between [0, 1]) using different precisions and splits. Accuracy is measured as the Frobenius norm of the difference between the double precision reference result and the Ozaki scheme result. From this data, we see one crucial point about the Ozaki scheme: the floating point precision for accumulation determines the overall precision. Hence the Ozaki scheme is particularly effective at taking advantage of hardware like NVIDIA’s Tensor Cores.
3 Implementation Details
@@
For this study, we created two different implementations of density matrix purification: one based on NVIDIA’s CUDA API for use on an NVIDIA RTX A6000 GPU and another using GNU MPFR to emulate low precision operations in software. MPFR will be particularly helpful to this study as it allows us to investigate arbitrarily defined floating point numbers 62. Thus we will use MPFR to explore precise precision requirements, and validate our results using the (substantially faster) CUDA implementation.
@@
As input to our purification implementation, we use Hamiltonians coming from the PySCF 63 and BigDFT 64 codes. We use HGH pseudopotentials 65 to remove the core electrons in BigDFT. BigDFT calculations are performed in the linear-scaling mode in order to produce a Hamiltonian in an LCAO basis set 66. PySCF calculations are done with the Polarization Consistent basis sets series 67. Calculations with BigDFT are performed with the PBE exchange-correlation functional 68 and for PySCF with B3LYP 69.
For simplicity, we first transform equation 4 to the standard eigenvalue problem using the Löwdin method:
$\ddot{H}=S^{-\frac{1}{2}}HS^{-\frac{1}{2}},$ (8)
$K=S^{-\frac{1}{2}}\ddot{K}S^{-\frac{1}{2}}.$ (9)
This transformation is performed by computing the inverse square root of the overlap matrix and transforming the Hamiltonian in double precision (though this could potentially be done with a diagonalization free method in low precision 48). All analysis of errors is then performed using the orthogonalized Hamiltonian and density matrix.
@@
For our tests, we implement the Trace Resetting Fourth Order (TRS4) purification method 40. We choose this method because it gives a clearer convergence signal than lower order methods. As a convergence criterion for the purification iterations, we use a change in the electronic energy, $Tr(\ddot{H}\ddot{K})$, of $1\times 10^{-8}$ Hartree. Due to the introduction of errors from simulating low precision, the purification algorithm may not be able to converge. For cases where convergence can’t be achieved (due to low precision errors), the purification algorithm halts when the previous two energy values have increased, with an absolute change of less than $1\times 10^{-2}$ Hartree (or on the 100th iteration). These failures usually happen when the final energy error is above $1\times 10^{-8}$ Hartree, though there are some exceptions. The source code of our implementation is available online (gitlab.com/wddawson/ozp), including Jupyter notebooks for running the experiments (with extra data about convergence detection and exponent underflow).
3.1 Error Evaluation Criteria
In this work, we will attempt to evaluate density matrix purification based on low precision arithmetic as a replacement for the dense eigenvalue solver used in LCAO codes. Of particular importance for such a replacement is that it should not impact convergence of the self-consistent field (SCF) iterations due to numerical noise. If a reliable ground state solution is found, properties may then be computed using a different precision. In Supplementary Information II, we demonstrate this for the case of the nuclear gradients.
To monitor SCF convergence, codes based on an LCAO basis set use a variety of different measures. In NWChem 7.2.2, convergence is based on three different criteria. First, there is the root mean square deviation (RMSD) of changes in the density matrix. Second, there is the change in the electronic energy, $Tr(\ddot{H}\ddot{K})$, between iterations. Third, the norm of the change in the commutator DIIS error vector is used 70:
$||\ddot{H}\ddot{K}-\ddot{K}\ddot{H}||_{F}.$ (10)
The default convergence criteria for these measures when performing DFT calculations are $1\times 10^{-5}$ (RMSD), $1\times 10^{-6}$ Hartree (energy), and $5\times 10^{-4}$ (DIIS), respectively. Some other codes use stricter convergence criteria by default. In Gaussian 16, the default cutoff for the RMSD is $1\times 10^{-8}$. In PySCF 2.7.0, the convergence is based on the change in the energy being below $1\times 10^{-9}$.
Here we target to achieve a precision in the RMSD, energy, and DIIS error vector at about two orders of magnitude below each of the default settings for NWChem. An error above this in the density matrix may make it difficult to test for convergence as numerical precision leads to fluctuations. An error in the energy may also lead to unacceptable noise in properties like energy differences. If the Hamiltonian does not commute with the density matrix, the optimization procedure will take steps related purely to numerical noise, which could severely hamper convergence. Approximately two orders of magnitude would also ensure that tighter convergence thresholds can be used when needed, such as when computing numerical gradients to build the Hessian matrix.
4 Numerical Experiments
@@
As test cases, we will first use the Hamiltonian coming from a BigDFT calculation of a Molnupiravir molecule bound to six water molecules. We will then expand the dataset to include a selenite ion surrounded by 10 water molecules calculated with different basis sets in PySCF. We also will include water clusters of different sizes and bulk silicon computed with BigDFT. We finally analyze a large water cluster (573 molecules) computed with B3LYP/PCSEG using NTChem 71, 72 (due to the system size). The main systems used in this paper are show in Fig 4.
4.1 Precision Requirements Sweep
@@
In order to understand the precision requirements for computing the density matrix, we perform density matrix purification using different custom floating point types implemented using MPFR on the Molnupiravir system. For this experiment, we assume double precision range for the exponent. In Fig. 5, we plot the error of the resulting density matrix, commutator, and energy compared to a reference result computed with diagonalization in double precision. We vary the effective precision of the mantissa used for multiplication from 11 (half) to 53 (double). For each calculation, we use one of three fixed accumulation mantissa values (24, 37, 53) or the “same” precision as multiplication.
@@
Figure 5: Errors when using purification vs. various mantissa sizes (including the implicit bit) for multiplication and accumulation applied to the Molnupiravir / BigDFT system (matrix dimension $147\times 147$). Different lines represent different values of accumulation precision and the horizontal axis represents the precision used for multiplication. The “same” line uses the same precision for multiplication and precision. The gray box extends from the NWChem convergence cutoffs down to two orders of magnitude lower. “Mixed Abs. Energy” labels the energy error after applying one McWeeny step in full double precision.
From this data, we see that single precision is not sufficient to achieve a converged result. The errors in the density matrix and commutator are not quite below the two orders of magnitude range set by our NWChem target (much less the codes with stricter criteria). Interestingly, we observed that the error in the energy was significantly worse compared to the other measures. Following the suggestion Finkelstein et al. 45, we investigated a mixed precision approach, where following the the convergence of the purification algorithm we perform one purification step using full double precision. Specifically, in our case we take one McWeeny step:
$P=3P^{2}-2P^{3}.$ (11)
We found that applying this iterative refinement substantially improved the energy values, though this improvement did not extend to other measures. For example, with a single precision mantissa the absolute energy error improved from $1.08\times 10^{-5}$ to $4.67\times 10^{-11}$; however, the RMSD and commutator norm only changed from $1.13\times 10^{-7}$ to $1.09\times 10^{-7}$ and $1.00\times 10^{-5}$ to $9.99\times 10^{-6}$, respectively. We further found that the magnitude of the error in the RMSD for any given multiplication in the iterations remains roughly the same, so there would be little benefit to adjusting the precision across iterations. Overall, we recommend the iterative refinement strategy when precision in the energy is important (such as the final SCF iteration).
@@
Another interesting observation is that a substantial amount of error in single precision comes from the accumulation phase, not the multiplication step. Importantly, we found that 37 bits of effective precision is sufficient for accumulation; for a single precision multiplication with 37 bits of accumulation, the errors are nearly indistinguishable from double precision accumulation. It is thus possible to obtain a high precision result while storing (and hence communicating) intermediate matrices in single precision if they are multiplied in a higher precision (a point we will return to in Sec. 4.6). Overall, we find that the double precision commonly used for quantum chemistry calculations essentially provides an unnecessarily high level of precision for computing the single particle density matrix.
@@
Figure 6: Errors when using purification vs. various mantissa sizes (including the implicit bit) for multiplication and accumulation applied to the selenite / PCSEG-3 system (matrix dimension $1508\times 1508$). Different lines represent different values of accumulation precision and the horizontal axis represents the precision used for multiplication. The “same” line uses the same precision for multiplication and precision. The gray box extends from the NWChem convergence cutoffs down to two orders of magnitude lower. “Mixed Abs. Energy” labels the energy error after applying one McWeeny step in full double precision.
We now repeat this experiment using the selenite system with the PCSEG-3 basis set (Fig. 6). The overall pattern is very similar, however the errors are shifted up compared to the BigDFT calculation. We thus conclude that the quadruple-$\zeta$ basis set represents more challenging numerical conditions and hence requires larger floating point data types. Even in this case though, FP64 provides far more precision than is required. If refinement is applied to the energy values, the 37 effective bit mantissa provides a sufficient amount of precision, and may provide acceptable results when used to accumulate the multiplication of single precision matrices. Overall, we find that low precision algorithms show promise for these problems, particularly if they can efficiently exploit the fact that less than full double precision is required.
4.2 Ozaki Scheme Application